Part A Selpat Compeany produoes silicons baking man in two processes heating and swnong An materials art iddd麒the begmng of de lleing Deprtncnt proces. The coneuny ssed weighood-averagc method io compute oqivaless usiis On Otober 1 Heating Departmeal precessed 2,000 mis tat were 70% oorplele. DunROcǐher, n complete were in procos Below is the cost imformalion Soe Healing Deportment. Wark in processi1 Octobor Conversion costs Total cost of work in process (1 Oetober Cot incurred in Oclobe Materials Labour .600 Required Prepare a peoduction cost repoet for Heating Department for mens of Ocober 2017 (1S marks) ITOTAL IS MARKS End ef Pag
Answer #1

Silpat Company

Production Cost Report( Heating Department)

For the month ended 31st October 2017

  1. Quantity Schedule

Units

Units

                      Units in process in the beginning

2,000

                       Units started & completed&transferred

                                                  To next department

20,000

                      Units in process at the end

4,000

26,000

  1. Cost Charged To The Department

                     Material( 1280+6000)

7280

                    Labour(720+4600)

5320

                    Overhead(2000+8100)

10100

22700

Cost Charged To The Department

22700

  1. Equivalent units produced

Material

23000

Labour and Overhead

23000

  1. Cost per unit

Material( 7280/23000)

0.317

Conversion Cost(15420/23000)

0.670

0.987

Cost per unit

0.987

  1. Cost Accounted for as follows

Opening Work-in-process

Material(600*0.317)

190.2

Conversion cost(600*0.670)

402

592.2

Units introduced and transferred( 20,000*0.987)

19740

Closing Work-in-process(2400*0.987)

2368.8

22701

Equivalent Units Produced

Materials {(2000*0.3)+(20,000*1.0)+(4000*0.6)}= 23000 units

Labour costs{(2000*0.3)+(20,000*1.0)+(4000*0.6)}=23000 units

Overhead costs{(2000*0.3)+(20,000*1.0)+(4000*0.6)}=23000 units

Unit cost of Equivalent unit for a given period is as follows:

Unit Cost= Costs of the period/Equivalent units of the period

Cost Per Unit

Materials(R7200/23000 units)=0.317

Overheads(RM